social class VS. number of children in the family is tabulated below: find the probabilities.


No of children lower middle upper

0 7 18 6

1 9 38 23

2 34 97 58

3 37 31 30


  1. P ( 3 children)
  2. P ( Lower or 2 children)
  3. P ( Middle and 1 child)
  4. P ( lower / 2 children)
  5. P ( 0 child / Middle)
1
Expert's answer
2021-02-24T07:10:45-0500

1.Total number of families:

"N=7+18+6+9+38+23+34+97+58+37+31+30=392"

P ( 3 children)"=\frac{37+31+31}{392}=0.25"


2.

P ( Lower or 2 children)"=\frac{7+9+34+37+97+58}{392}=0.62"


3.

P ( Middle and 1 child)"=38\/392=0.097"


4.

P ( lower / 2 children)"=\frac{7+9+37}{392}=0.135"


5.

P ( 0 child / Middle)"=\frac{7+6}{392}=0.033"